Please help me :(
2(x-3)=-6
Thank you.

Answer by uma(370)

2(x - 3) = - 6
Removing the parenthesis by multiplying the outer value with each of the term inside:
2x - 6 = -6
==> 2x - 6 + 6 = -6 + 6 [Adding 6 to both the sides]
==> 2x = 0
==> 2x/2 = 0/2
==> x = 0
x = 0 is the solution for the given equation.
Good Luck!!!